OpenStudy (anonymous):

Help me please! I will give a medal :) solve the equation (3x+2)(x-5)=0 enter the solution with the highest value

rvc (rvc):

multiply it

rvc (rvc):

do u know?

rvc (rvc):

@jad2015 what happened?

OpenStudy (anonymous):

@rvc yes I know how to multiply it, but shouldn't the highest value be 5

rvc (rvc):

\[(a+b)(c+d)=a(c+d)+b(c+d)\]

rvc (rvc):

so what u get?

rvc (rvc):

\[3x(x-5)+2(x-5)=3x^2-15x+2x-10\]

rvc (rvc):

so what u get now? @jad2015

rvc (rvc):

please reply

OpenStudy (anonymous):

@rvc when multiplied the outcome is 3x^2-13x-10

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

i hope u got it @jad2015

OpenStudy (anonymous):

if not plaese say